OpenStudy (calculusxy):

Solve: 3 1/3 X 3 3/4 X 2/5=

OpenStudy (calculusxy):

@shamil98

OpenStudy (shamil98):

10/3 x 15/4 x 2/5 Think of this as. \[\large \frac{ 10 \times 15 \times 2 }{ 3 \times 4 \times 5 }\]

OpenStudy (calculusxy):

Oh. Made a silly mistake. I replaced a 3 instead of 4 in the third fraction. :/ And the answer is 5. Thanks for your help. But can you help with another one?

OpenStudy (calculusxy):

*second fraction

OpenStudy (shamil98):

sure.

OpenStudy (calculusxy):

If the 5% sales tax on a snowmobile was $42, what was the price of the snowmobile not including tax? A. $840 B. $210 C. $820 D. $640

OpenStudy (shamil98):

Basically, if 5% of something is 42 how much is that something? 0.05x = 42 solve for x.

OpenStudy (calculusxy):

You have to divide 42 by 0.05 right?

OpenStudy (calculusxy):

Would that be $840?

OpenStudy (shamil98):

Yes

OpenStudy (calculusxy):

Mr. Summon paid $58.50 in interest on a loan that had 6% interest rate. How much did he borrow?

OpenStudy (calculusxy):

Is it $975? I got it cuz I did 58.50/0.06=975

OpenStudy (shamil98):

He paid 58.50 of interest, just subtract 975-58.50 =

OpenStudy (calculusxy):

That results to 91.75 and the answer choice doesn't have that.